# Universal Approximation Theory: Foundations for Parallelism in Neural Networks

### v4

@@ 3.2 Para-former
3.3 Speed-up Ratio
We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks. It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging. Therefore, we will focus on a theoretical analysis. Based on the network structure shown in Figure 3 and 4, the inference time of Para-Former is only affected by its network depth.
4 Experiments
Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the UAT approximation properties (The larger the value of $N$, the more powerful the model becomes, resulting in better classification performance.) using several commonly used public image classification datasets (Section 4.2). We then explore the impact of network depth on the Para-Former performance (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4). We conducted experiments with various Para-Former models, where the naming convention uses the first number to indicate the depth and the second number to indicate the number of layers. For instance, Para-former-1-12 denotes a model with 1 depth and 12 layers.
4.1 Dataset
In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis. These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models. Specifically, our study centers on the following pivotal datasets: CIFAR-10, CIFAR-100 (Krizhevsky 2009), STL-10 (Coates, Ng, and Lee 2011), Flowers-102 (Nilsback and Zisserman 2008) and Oxford-IIIT Pets (Parkhi et al. 2012). In Table 1, we provide more details for these datasets.
@@ 4.1 Dataset
4.2 The Feasibility of Para-Former
To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs. The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding. We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figure 3 (We evaluated the models’ performance using the accuracy metric, calculated as $\text{accuracy}=100\times\frac{\text{correct}}{\text{total}}$). A single-layer setup corresponds to a ViT with a depth of 1. The results of the models are presented in Table 2.
It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.). According to UAT theory, increasing the network layers is equivalent to increasing $N$ in Eq. (1), thereby enhancing the approximation capability of UAT. The results in Table 2 are in complete agreement with UAT theory. Therefore, we believe that this parallel network approach is entirely feasible.
@@ 4.2 The Feasibility of Para-Former
1. Model Fitting Capability: This is influenced by the model’s size and design, particularly the value of $N$ in the UAT and the degrees of freedom of the parameters. A larger $N$ enhances the UAT’s ability to fit functions. The degrees of freedom refer to the number of parameters in the UAT that are influenced by the input. For instance, the degrees of freedom in Equation (8) is 2, while in Equation (9), it is 6 (see Appendix B). The degrees of freedom represent the model’s ability to dynamically fit corresponding functions based on the input. Simply increasing $N$ can improve the model’s ability to fit specific functions, but images are usually diverse. Thus, the network must have the capability to dynamically fit different functions based on the input. Higher degrees of freedom allow the model to fit multiple different functions according to the input. Additionally, the degrees of freedom vary within the Transformer: they mainly consist of weights and biases. The degrees of freedom for weights stem from the MHA, primarily putting linear transformations based on the input. The degrees of freedom for biases come from residual connections, which can dynamically fit the bias terms using UAT. The deeper the residual layers, the more UAT layers correspond to the bias terms, leading to stronger dynamic fitting capabilities. As a result, a shallow serial network can achieve the fitting ability of a multi-layer parallel network with a depth of 1. We will explore this issue further in Section 4.3.
@@
Figure 5: OxfordIIITPet dataset’s accuracy on different classes.
Figure 6: Flower-102 dataset’s accuracy on different classes.
@@ 4.3 Model Depth Effect
Based on the conclusions from Section 4.2, this section explores the impact of network depth on prediction performance. According to equations (8) and (9), the depth of the model primarily affects the degrees of freedom of the model parameters within the UAT. Therefore, we designed different versions of the Para-Former model with 24 layers and depths of 2, 3, and 6, comparing their predictive performance against a ViT with a depth of 6. The results are shown in Table 3. The degrees of freedom for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are 48, 18, 72, 72, and 72, respectively (see Appendix B).
Why does Para-F-6-6, despite having the lowest degrees of freedom, only perform worse than Para-F-6-24? This is because the degrees of freedom encompass both the freedom of weights and biases. The freedom of weights can be understood as linear transformations derived from inputs, whereas the freedom of biases results from the approximation by multiple layers of UAT, which allows for stronger dynamic fitting capabilities. The deeper the layers, the stronger the approximation ability. Thus, Para-F-6-24 performs the best, followed by Para-F-6-6, as the number of UAT layer of their bias terms is largest (see Appendix B).
@@ 4.4 Dataset Effect
Figure 7: The fine-tuning process from the perspective of UAT.
@@
Figures 5 and 6 show some results from the OxfordIIITPet and Flower-102 datasets. Both figures indicate that for categories with diverse backgrounds and inherent feature diversity, the prediction results are poor, as shown in Figures 5.b and 6.b. In contrast, when the background is consistent but the object morphology is diverse, the results are relatively better, as shown in Figures 5.a and 6.a. The best results are achieved when both the background and object morphology are minimal, as shown in Figures 5.c and 6.c. Although the Flower-102 dataset has fewer images, its backgrounds are more consistent, and the differences between species are smaller. Conversely, animals in the OxfordIIITPet dataset can appear in various forms, whereas flowers in the Flower-102 dataset are usually photographed from relatively fixed angles, leading to better performance.
Does this mean that deep learning models are not powerful enough or lack generalization and reasoning capabilities? Quite the opposite. With such limited data, deep learning models achieved about 25% accuracy on OxfordIIITPet and about 40% on Flower-102, actually demonstrating the strength and generalization ability of deep learning models. A common misconception about deep learning is that even if people are unfamiliar with certain species, by careful examination, they can accurately distinguish between OxfordIIITPet and Flower-102 images, leading to the belief that deep learning lacks generalization or human-like reasoning capabilities. This view is flawed. Suppose the human brain operates like a UAT model; it can automatically ignore background information before seeing these images because such backgrounds are frequently encountered in daily life, and the UAT in the brain has already fitted this information. Therefore, people only need to focus on specific information and then use this information to train the brain to identify similarities and differences. For example, a 10-year-old child’s brain, observing the world at 10 frames per second (Read and Meyer 2000), would have processed over $1.5\times 10^{9}$ high-resolution images by age 10 (Lu et al. 2017), many containing common objects and backgrounds from daily life. Hence, it is unfair to criticize deep learning models for being unable to train on small datasets and for lacking universality and generalization ability.
@@ 5 Conclusion
In this paper, we propose a parallel network based on the UAT and validate its effectiveness on large datasets. We analyze the characteristics of both serial and parallel networks. Serial networks based on residual connections, often have higher degrees of parameter freedom, leading to stronger generalization capabilities. However, as the network depth increases, the inference time also becomes longer. On the other hand, parallel networks trade computational resources for time, meaning their runtime is not affected by the number of network layers. This allows for the design of multi-layer parallel networks that can replace serial networks to achieve fast inference. Additionally, we analyze the impact and reasons data have on deep learning. From the UAT perspective, training on large datasets essentially enhances the model’s ability to fit diverse data.
Appendix A The UAT Format of Transformer-based ViTs
In this section, we will demonstrate that the mathematical form of single-layer and two-layer Transformers is equivalent to the UAT. Firstly, the basic form of a Transformer is shown in Figure 2 of main text. We represent the MHA and FFN in the Transformer in matrix-vector form as Eq. (10) and Eq. (11), respectively (for specific details, refer to UAT2LLMs). To distinguish between the operations and variables in the Transformer and those in the matrix-vector form, we add a prime () symbol to the operations and variables in the matrix-vector form. It is important to emphasize that the parameters of the MHA are influenced by the input.
$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (10)
$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (11)
@@ Appendix A The UAT Format of Transformer-based ViTs
Let $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, so we obtain:
$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (13)
where $\mathbf{W}^{\prime}_{1,1}$ and $\mathbf{W}^{\prime}_{1,2}$ are parameters influenced by the input, thus having a degree of freedom of 2.
$\mathbf{x}^{\prime}_{2}$ $=\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})+\mathbf{b}^{\prime}_{2,3}$ (14)
$=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,3})+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})$
=$=\{\mathbf{W}^{\prime}_{2,1}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,3}\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,2}\}$
=$=\{\mathbf{W}^{\prime}_{2,1}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,3}+\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}({\mathbf{W}}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,2}+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
=$=\{(\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}})+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}\}$
In Eq. (14), we provide the matrix-vector format of 2-layer Transformer. Let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}$, $\mathbf{b}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$. Thus, Eq. (14) can be written as:
$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (15)
$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
where $\mathbf{W}^{\prime}_{2,1}$,$\mathbf{b}^{\prime}_{2,1}$,$\mathbf{W}^{\prime}_{1,2}$,$\mathbf{W}^{\prime}_{1,3}$,$\mathbf{W}^{\prime}_{2,3}$, and $\mathbf{b}^{\prime}_{2,2}$ are influenced by the input, so the degree of freedom is 6.
Another issue is that in $\mathbf{x}^{\prime}_{1}$ and $\mathbf{x}^{\prime}_{2}$, there is a term $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ and $(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})$ which doesn’t generally conform to the standard terms in UAT. We can consider this term as a special case and regard it as a layer in UAT. Firstly, when the number of network layers increases, it effectively increases $N$ in UAT. Given UAT’s strong approximation capabilities, adding multiple instances of this term can be viewed as a linear transformation, thus it doesn’t affect the overall approximation power of UAT. Furthermore, this term can be considered a special form in UAT. Taking ReLU as an example activation function, $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ can be expressed as $\text{ReLU}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})-\text{ReLU}(-(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3}))$.
In summary, we have shown that the mathematical forms of one-layer and two-layer Transformers can be seen as 2 and three-layer UAT.

### v5

@@ 4.1 Dataset
4.2 The Feasibility of Para-Former
To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs. The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding. We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figures 4 and 5 (We evaluated the models’ performance using the accuracy metric, calculated as $\text{accuracy}=100\times\frac{\text{correct}}{\text{total}}$). A single-layer setup corresponds to a ViT with a depth of 1. The results of the models are presented in Table 2.
It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.). According to DUAT theory, increasing the network layers is equivalent to increasing $N$ in Eq. equation 1, thereby enhancing the approximation capability of DUAT. The results in Table 2 are in complete agreement with DUAT. Therefore, we believe that this parallel network approach is entirely feasible.
@@ 4.2 The Feasibility of Para-Former
1. Model Fitting Capability: This is influenced by the model’s size and design, particularly the value of $N$ in the DUAT and the degrees of freedom of the parameters. A larger $N$ enhances the DUAT’s ability to fit functions. The degrees of freedom refer to the number of parameters in the DUAT that are influenced by the input. For instance, the degrees of freedom in Eq. equation 5 is 2 (see Appendix B). The degrees of freedom represent the model’s ability to dynamically fit corresponding functions based on the input. Simply increasing $N$ can improve the model’s ability to fit specific functions, but images are usually diverse. Thus, the network must have the capability to dynamically fit different functions based on the input. Higher degrees of freedom allow the model to fit multiple different functions according to the input. Additionally, the degrees of freedom vary within the Transformer: they mainly consist of weights and biases. The degrees of freedom for weights stem from the MHA, which can be simply viewed as a generated parameters based the covariance matrix of the input. The degrees of freedom for biases come from residual connections, which can dynamically fit the bias terms using DUAT. The deeper the residual layers, the more DUAT layers correspond to the bias terms, leading to stronger dynamic fitting capabilities. As a result, a shallow serial network can achieve the fitting ability of a multi-layer parallel network with a depth of 1. We will explore this issue further in Section 4.3.
@@ 4.3 Model Depth Effect
Based on the conclusions from Section 4.2, this section explores the impact of network depth on prediction performance. According to equations equation 5 and equation 6, the depth of the model primarily affects the degrees of freedom of the model parameters within the DUAT. Therefore, we designed different versions of the Para-Former model with 24 layers and depths of 2, 3, and 6, comparing their predictive performance against a ViT with a depth of 6. The results are shown in Table 3. The degrees of freedom for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are 48, 15, 72, 72, and 60, respectively (see Appendix B).
Why does Para-F-6-6, despite having the lowest degrees of freedom, only perform worse than Para-F-6-24? This is because the degrees of freedom encompass both the freedom of weights and biases. The freedom of weights can be understood as linear transformations derived from inputs covariance, whereas the freedom of biases results from the approximation by multiple layers of DUAT, which allows for stronger dynamic fitting capabilities. The deeper the layers, the larger number of $N$ in DUAT for bias term, and the stronger the approximation ability. Thus, Para-F-6-24 performs the best, followed by Para-F-6-6, as the number of DUAT layer of their bias terms is largest (see Appendix B).
@@ 4.4 Dataset Effect
Figure 6: OxfordIIITPet dataset’s accuracy on different classes.
Figure 7: Flower-102 dataset’s accuracy on different classes.
Figures 6 and 7 show some results from the OxfordIIITPet and Flower-102 datasets. Both figures indicate that for categories with diverse backgrounds and inherent feature diversity, the prediction results are poor, as shown in Figures 6.b and 7.b. In contrast, when the background is consistent but the object morphology is diverse, the results are relatively better, as shown in Figures 6.a and 7.a. The best results are achieved when both the background and object morphology are minimal, as shown in Figures 6.c and 7.c. Although the Flower-102 dataset has fewer images, its backgrounds are more consistent, and the differences between species are smaller. Conversely, animals in the OxfordIIITPet dataset can appear in various forms, whereas flowers in the Flower-102 dataset are usually photographed from relatively fixed angles, leading to better performance.
Figure 8: The fine-tuning process from the perspective of UAT.
Does this mean that deep learning models are not powerful enough or lack generalization and reasoning capabilities? Quite the opposite. With such limited data, deep learning models achieved about 25% accuracy on OxfordIIITPet and about 40% on Flower-102, actually demonstrating the strength and generalization ability of deep learning models. A common misconception about deep learning is that even if people are unfamiliar with certain species, by careful examination, they can accurately distinguish between OxfordIIITPet and Flower-102 images, leading to the belief that deep learning lacks generalization or human-like reasoning capabilities. This view is flawed. Suppose the human brain operates like a DUAT model; it can automatically ignore background information before seeing these images because such backgrounds are frequently encountered in daily life, and the DUAT in the brain has already fitted this information. Therefore, people only need to focus on specific information and then use this information to train the brain to identify similarities and differences. For example, a 10-year-old child’s brain, observing the world at 10 frames per second Read & Meyer (2000), would have processed over $1.5\times 10^{9}$ high-resolution images by age 10 Lu et al. (2017), many containing common objects and backgrounds from daily life. Hence, it is unfair to criticize deep learning models for being unable to train on small datasets and for lacking universality and generalization ability.
@@ 5 Conclusion
In this paper, we propose design guidelines for parallel neural networks based on DUAT, which, to our knowledge, is the first work to establish principles for parallel networks. Based on this theory, we designed a parallel network, Para-Former, and validated its effectiveness across several commonly used datasets. Through DUAT, we analyzed the characteristics of both serial and parallel networks. Serial networks with residual connections often have higher parameter flexibility, leading to better generalization capabilities. However, as network depth increases, inference time also grows. In contrast, parallel networks trade off computational resources for time efficiency, meaning their runtime is unaffected by the number of layers. This allows us to design multi-layer parallel networks that can replace serial networks, achieving faster inference. Additionally, we examined the impact of data on deep learning performance and explored the underlying reasons. From the DUAT perspective, training with large datasets essentially enhances a model’s capacity to fit diverse data.
Appendix A The DUAT Format of single layer ViT
In this section, we will demonstrate that the DUAT form of a single-layer layer ViT . Firstly, the basic form of a Transformer is shown in Figure 2 of main text. We represent the MHA and FFN in the Transformer in matrix-vector form as Eq. equation 8 and Eq. equation 9, respectively (for specific details, refer to DUAT2LLMs). To distinguish between the operations and variables in the Transformer and those in the matrix-vector form, we add a prime () symbol to the operations and variables in the matrix-vector form. It is important to emphasize that the parameters of the MHA are influenced by the input.
$MHA^{\prime}(\mathbf{x}^{\prime}_{i})\mapsto\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (8)
$FFN^{\prime}(\mathbf{x}^{\prime}_{i})\mapsto\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (9)